OpenStudy (hannahlb99):

Solving inequalities... help! (x-10)^3 > 53 A. x < -13.756 or x > 13.756 B. x < -7.28 or x > 7.28 C. x > 13.756 D. x > 15.418

alones (alones):

Alright so let's find the critical points of the inequality \[\ x^3−30x^2+300x−1000=53\]

alones (alones):

So subtract 53 from both sides \[\ x^3−30x^2+300x−1000−\color{red}{53}=53−\color{red}{53}\] \[\ x^3−30x^2+300x−1053=0\]

alones (alones):

I believe you know what cubic formula is use it \(x=13.756286\) Also don't forget to check intervals in between critical points.

alones (alones):

and your final answer should be \(x>13.756286\) since that would work in original inequality.

alones (alones):

@hannahlb99 when you get online tell me if you understand this.

OpenStudy (hannahlb99):

okay! thank you so much

OpenStudy (sshayer):

cube of a negative number is negative and cube of a positive number is positive. as right hand side is positive so x-10>0 x>10 \[x-10>53^{\frac{ 1 }{ 3 }}\] \[x-10>3.75628575\] x>10+3.75628575 ?